OpenStudy (anonymous):

I will give award a medal to whoever helps me with this question: find the area for each circle with a radius of 9 in.

OpenStudy (anonymous):

The formula of finding the area of a circle is: Area = Pi x r2 The value of Pi is always taken as 3.14. We are given the radius in the question, which is 9. To find the area, we will simply put in these values in the equation: Area = 3.14 x (9)2 Area = 3.14 x 81 = 254.34 square inches So, the area of the circle is 254.34 square inches.

OpenStudy (texaschic101):

area of a circle is : A = (pi)(r^2) A = 3.14(9^2) A = 3.14(81) A = 254.34

OpenStudy (anonymous):

I wish I could give u both medals

OpenStudy (anonymous):

Thats okay. I can help with that. I'll give texaschic101 a medal for you Punkchick! It'd be my pleasure! xD

OpenStudy (texaschic101):

thank you JoMary :)

OpenStudy (anonymous):

It was my pleasure texaschic101! :)